# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v3

@@ Abstract
Recent findings by Cohen et al. (2021) demonstrate that when training neural networks with full-batch gradient descent with step size $\eta$, the largest eigenvalue $\lambda_{\max}$ of the full-batch Hessian consistently stabilizes at $\lambda_{\max}=2/\eta$. These results have significant implications for convergence and generalization. This, however, is not the case of mini-batch stochastic gradient descent (SGD), limiting the broader applicability of its consequences. We show that SGD trains in a different regime we term Edge of Stochastic Stability (EoSS). In this regime, what stabilizes at $2/\eta$ is Batch Sharpness: the expected directional curvature of mini-batch Hessians along their corresponding stochastic gradients. As a consequence, $\lambda_{\max}$–which is generally smaller than Batch Sharpness–is suppressed, aligning with the long-standing empirical observation that smaller batches and larger step sizes favor flatter minima. We further discuss implications for mathematical modeling of SGD trajectories.
Edge of Stability, Optimization for deep learning, SGD
@@ 1 Introduction
Training algorithms are a key ingredient to the success of deep learning. Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951), a stochastic variant of Gradient Descent (GD), has been effective in finding parameters that yield good test performance despite the complicated nonlinear nature of neural networks.
In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2019, 2020; Cohen et al., 2021, 2022). Precisely, if the step size is $\eta>0$ for full-batch GD, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations. Cohen et al. (2021) named this regime Edge of Stability (EoS). Surprisingly, unlike for quadratics, where this would imply divergence, this instability does not damage convergence when training neural networks.
@@
The picture for full-batch algorithms seems (empirically) clear, both for GD (Cohen et al., 2021; Damian et al., 2023) and its adaptive and accelerated version (Cohen et al., 2022, 2024). Unfortunately, neural networks are typically trained with the mini-batch versions of these gradient-based methods, and how Cohen et al. (2021) in Section 6 and Appendices G and H noticed and stated in the introduction and limitations of their work, what they observed is not the case of mini-batch training. Precisely, to quote Cohen et al. (2021):
@@
while the sharpness does not flatline at any value during SGD (as it does during gradient descent), the trajectory of the sharpness is heavily influenced by the step size and batch size (Jastrzębski et al., 2019, 2020), which cannot be explained by existing optimization theory. Indeed, there are indications that the “Edge of Stability” intuition might generalize somehow to SGD, just in a way that does not center around the sharpness. []delimited-[][$\ldots$][ … ] In extending these findings to SGD, the question arises of how to model “stability” of SGD .
We define here a notion of mini-batch sharpness which acclimates to the learning rate and batch size in a way that mirrors $\lambda_{\max}$ in the Edge of Stability.
What was empirically known for mini-batch SGD .
In the case of mini-batch algorithms, (i) Jastrzębski et al. (2019, 2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when. Moreover, Cohen et al. (2021) noticed that SGD somehow acclimates to the hyperparameters. However, they did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 1, often without a proper progressive sharpening phase. This scenario leaves the most basic questions open: In what way the location of convergence of SGD acclimates to the choice of hyperparameters? What are the key quantities involved? To be more specific, can we characterize the training phenomena in (i), (ii), (iii) above? What determines them? Does SGD train in an unstable regime?
Edge of Stochastic Stability.
We address the questions above by introducing a general framework for SGD dynamics through the concept of Batch Sharpness: the expectation of the Hessians of mini-batch losses along the direction of the mini-batch gradients, see Section 3.1. This serves as a generalized and direct substitute for $\lambda_{\max}$ in the mini-batch SGD setting. Specifically, we identify a new regime–which we term Edge of Stochastic Stability (EoSS)–where Batch Sharpness stabilizes near $2/\eta$, while maintaining a critical gap with $\lambda_{\max}$. This dynamic explains the plateaus of $\lambda_{\max}$ at lower levels determined by batch size and learning rate, unifying the observed phenomena in mini-batch training.
Organization of the rest of the paper.
Section 2 reviews related work and outlines key open questions. In Section 3, we introduce and empirically validate the phenomenon of Edge of Stochastic Stability. Section 4 discusses what the observed notion of stability entails mathematically. In Section 5, we examine stability mathematically and we study values at which $\lambda_{\max}$ plateaus. In Section 6, we discuss the implications for the structure and modeling of SGD dynamics. Finally, Section 7 summarizes our findings and addresses the limitations of this work.
@@
Figure 1: Mini-batch SGD on CIFAR-10. The full-batch Hessian’s largest eigenvalue $\lambda_{\max}$ typically plateaus below $2/\eta$, here $\eta=1/200$. Smaller batch sizes lead to lower plateau values.
Figure 2: SGD at EoSS under different learning rates and batch sizes. We train a fully-connected network on an 8k subset of CIFAR-10. Batch Sharpness stabilizes at the $2/\eta$ threshold across varying batch sizes and step sizes.
2 Related Work
2.1 Progressive Sharpening and Edge of Stability
Progressive sharpening.
Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018, 2019; Fort & Ganguli, 2019). Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training (Keskar et al., 2016; Sagun et al., 2016; Fort & Scherlis, 2019; Jastrzębski et al., 2019). Later, Jastrzębski et al. (2019, 2020) and Cohen et al. (2021) managed to make this precise, noticing that, along the trajectories of SGD and GD, $\lambda_{\max}$ increases steadily, often after a small number of steps of decrease. This phenomenon was called progressive sharpening by Cohen et al. (2021).
A phase transition.
Early studies (Goodfellow et al., 2016; Li et al., 2019; Jiang et al., 2019; Lewkowycz et al., 2020) revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction. Jastrzębski et al. (2020) attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening. This transition slows the convergence by confining the dynamics to a region of lower sharpness. Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape. Jastrzębski et al. (2019, 2020); Cohen et al. (2021, 2022) indeed showed that the phase transition comes at different points for different algorithms on the same landscapes. Jastrzębski et al. (2019, 2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it. Cohen et al. (2021, 2022) later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms.
Full-batch edge of stability.
After the phase transition above, GD and full-batch Adam train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value. The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics. Observations from Cohen et al. (2021, 2022) indicate that most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution. Lee & Jang (2023) explained why $\lambda_{\max}$ in this regime is generally above $2/\eta$, rarely aligning exactly. The reasons include: (i) nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and (ii) the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
EoS and convergence.
There is a growing body of work analyzing the surprising mechanism of EoS in training dynamics with GD. Classically, when gradients depend linearly on parameters, divergence occurs locally if $\eta>\frac{2}{\lambda}$, as demonstrated with one-dimensional parabolas (Cohen et al., 2021). Yet, neural networks often converge even when $\eta\geq\frac{2}{\lambda}$, likely due to the problem’s non-standard geometry. Damian et al. (2023) proposed an explanation under some, empirically tested, assumptions of alignment of third derivatives and gradients. Wang et al. (2022) showed that the product-like structure of shallow networks guarantees convergence with learning rates exceeding the EoS threshold. Beneventano & Woodworth (2025) showed that although convergence is guaranteed, it may be extremely slow, this may, however, benefit implicit regularization effects. For further exploration of convergence and implicit regularizarion in the EoS regime, we direct the reader to Arora et al. (2022); Ahn et al. (2022, 2023); Zhu et al. (2023); Lyu et al. (2023).
2.2 SGD, Hessian, and Generalization
SGD finding flatter minima.
Our result is inherently a result about mini-batch training improving flatness. Specifically, we explain why: {mdframed}
Training with smaller batches constraints the dynamics in areas with smaller eigenvalues for the full-batch Hessian.
This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016). Analogously, recent work strongly aligns with our result: Jastrzębski et al. (2021) demonstrates that large learning rates in SGD act similarly to penalizing the trace of the Fisher matrix–which approximates the Hessian during training and shares its top eigenspaces (Jastrzębski et al., 2018; Martens, 2020; Thomas et al., 2020)–in image classification tasks.
Sharpness and Generalization.
SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance (LeCun et al., 2012; Keskar et al., 2016; Goyal et al., 2017; Jastrzębski et al., 2018; Masters & Luschi, 2018; Smith et al., 2021). This advantage has been widely attributed to the flatness of the minima (Murata et al., 1994; Hochreiter & Schmidhuber, 1994; Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020). Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks (Izmailov et al., 2019; Foret et al., 2021). Moreover, Jastrzębski et al. (2021) shows that, in practice, penalizing the trace of the Fisher information matrix empirically improves generalization and reduces memorization—especially when training with noisy labels. We believe that our quantification and the introduction of a theoretical framework for these improvements in sharpness may be a stepping stone towards understanding these generalization benefits theoretically.

### v4

@@ Abstract
Recent findings by Cohen et al. [1] demonstrate that when training neural networks with full-batch gradient descent with step size $\eta$, the largest eigenvalue $\lambda_{\max}$ of the full-batch Hessian consistently stabilizes at $\lambda_{\max}=2/\eta$. These results have significant implications for convergence and generalization. This, however, is not the case of mini-batch stochastic gradient descent (SGD), limiting the broader applicability of its consequences. We show that SGD trains in a different regime we term Edge of Stochastic Stability (EoSS). In this regime, what stabilizes at $2/\eta$ is Batch Sharpness: the expected directional curvature of mini-batch Hessians along their corresponding stochastic gradients. As a consequence, $\lambda_{\max}$—which is generally smaller than Batch Sharpness—is suppressed, aligning with the long-standing empirical observation that smaller batches and larger step sizes favor flatter minima. We further discuss implications for mathematical modeling of SGD trajectories.
@@ 1 Introduction
The choice of training algorithm is a key ingredient in the deep learning recipe. Extensive evidence, e.g. [2], indeed shows that performance consistently depends on the optimizer and hyper-parameters. How and why this is the case is a key question in optimization theory and deep learning.
Gradient Descent (GD) was shown to optimize neural networks in a regime of instability, termed Edge of Stability (EoS) by Cohen et al. [1]. With a constant step size $\eta$, the highest eigenvalue of the Hessian of the full-batch loss—denoted here as $\lambda_{\max}$—grows until $2/\eta$ and hovers right above that value, subject to small oscillations [1, 3, 4, 5, 6]. Although classical convex optimization theory would call this step size “too large”, the loss continues to decrease. These works established a number of surprising facts: (1) that we require an optimization theory which works in more general scenarios then the classical $\eta<2/L$; (2) that GD for neural networks may not stabilize at stationary points but when $\lambda_{\max}=2/\eta$; (3) what a source of instability of (pre-)training is: progressive sharpening; (4) how location of stabilization depends on the choice of hyperparameters of full-batch optimizer.
While real-world training is almost always mini-batch—given the large amount of data—existing EoS analyses explicitly do not apply to this case: no curvature-type quantities, as $\lambda_{\max}$, are known to similarly affect SGD while training neural networks. We bridge this gap by establishing that mini-batch SGD trains in a similar regime of instability which we term Edge of Stochastic Stability (EoSS). Precisely, a quantity we call Batch Sharpness, see Definition 1 and Figure 1, hovers around $2/\eta$ and implicitly functions as sharpness in the case of mini-batch algorithms.
Organization and Contributions.
Section 2 reviews related work and outlines the key open questions we tackle, namely how does the Edge of Stability phenomenon generalize to mini-batch algorithms and how can we quantify the characters in play. As oscillatory behaviors are central to these phenomena, Section 3 disentangles the two types of SGD oscillations into noise-driven (as in Robbins-Monro–type of stochastic optimization when the step size is kept fixed) and curvature-driven—which are the ones we are interested in. We establish here that most training happens under oscillations, only the second part of those are curvature-driven. In Section 4, we introduce, properly characterize, and empirically validate the phenomenon of Edge of Stochastic Stability. In Section 5 we discuss the theoretical motivation behind the quantity we term as Batch Sharpness. Batch Sharpness is the implicit notion of curvature that governs the phenomenon of Edge of Stability in the case of mini-batch algorithms, see Definition 1. It is the expected curvature along the step on the mini-batch landscape. We also prove that when Batch Sharpness bigger than $2/\eta$, the oscillations are unstable. A good feature of full-batch–EoS phenomenon was that the quantity of interest (the largest eigenvalue of the Hessian) has a key geometrical role which has been linked both to generalization and to speed of optimization. For Batch Sharpness this link is less direct, we thus investigate in Section 6 the relationship between these two quantities. Our results are another proof of the fact that the dynamics of noise-injected GD or SDEs and the dynamics of mini-batch SGD are qualitatively different and studying the firsts could be misleading for inducing properties of the second. We discuss this implication in Section 7, before concluding in Section 8.
Figure 1: SGD at EoSS under different learning rates and batch sizes. MLP on an 8k subset of CIFAR-10 with step size $\eta>0$. Batch Sharpness stabilizes at the $2/\eta$ threshold across varying batch sizes and step sizes.
2 Related Work
Progressive sharpening.
Early studies observed that the local shape of the loss landscape changes rapidly at the beginning of the training [2, 4, 7, 8, 9, 10]. Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training [2, 4, 11, 12]. Subsequently, Jastrzębski et al. [4, 5] and Cohen et al. [1] precisely characterized this behavior, demonstrating a steady rise in $\lambda_{\max}$ along GD and SGD trajectories, typically following a brief initial decline. This phenomenon was termed progressive sharpening by Cohen et al. [1].
A phase transition.
Early works [13, 14, 15, 16] revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction. Jastrzębski et al. [5] attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening. This transition slows the convergence by confining the dynamics to a region of lower sharpness. Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape. Jastrzębski et al. [4, 5], Cohen et al. [1, 3] indeed showed that the phase transition comes at different points for different algorithms on the same landscapes. Jastrzębski et al. [4, 5] showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it. Cohen et al. [1, 3] later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms. We manage here to quantify the value of the instability threshold for mini-batch SGD, thus characterizing when the phase transition happen for SGD. Importantly, the fact that progressive sharpening does not end on the landscape, but is limited by algorithmic properties implies that the locations of stabilization of the optimization algorithms surprisingly are not stationary points.
Full-batch edge of stability.
After the phase transition above, GD and full-batch Adam train in the EoS oscillatory regime [1, 3, 6], where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value. The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics. Observations from Cohen et al. [1, 3] indicate that, under mean square error (MSE), most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution. Lee and Jang [17] explained why in this regime $\lambda_{\max}$ often slightly exceeds $2/\eta$: this deviation arises primarily from nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
EoS and convergence.
A growing body of research analyzes the surprising mechanism underlying EoS dynamics observed during training with GD. Classically, when gradients depend linearly on parameters, divergence occurs locally if $\eta>\frac{2}{\lambda_{\max}}$, as illustrated by one-dimensional quadratic models [1]. In contrast, neural networks often converge despite violating this classical stability condition, presumably due to the problem’s non-standard geometry. Damian et al. [18] propose an explanation under some, empirically tested, assumptions of alignment of third derivatives and gradients. For further insights into convergence and implicit regularization phenomena in the EoS regime, we direct the reader to [19, 20, 21, 22, 23, 24, 25, 26].
@@ 2 Related Work
While the empirical behavior of EoS for full-batch algorithms is relatively well-understood, neural networks are predominantly trained using mini-batch methods. As explicitly noted by Cohen et al. [1, Section 6, Appendices G and H], their observations and analysis do not directly apply to mini-batch training. In particular, they emphasize:
@@
[…] while the sharpness does not flatline at any value during SGD (as it does during gradient descent), the trajectory of the sharpness is heavily influenced by the step size and batch size Jastrzębski et al. [4, 5], which cannot be explained by existing optimization theory. Indeed, there are indications that the “Edge of Stability” intuition might generalize somehow to SGD, just in a way that does not center around the (full-batch) sharpness. […] In extending these findings to SGD, the question arises of how to model “stability” of SGD.
In this work, we show that the EoS phenomenon does indeed generalize to SGD, and we identify the central quantity governing this generalization (Batch Sharpness in Definition 1). We modeled stability of SGD on the neural networks landscapes: our answer is that empirically SGD is stable if on average the step is stable on the mini-batch landscape—not on the full-batch landscape.
What was empirically known for SGD.
@@
Figure 2: SGD on CIFAR-10: $\eta=1/200$. The full-batch Hessian’s $\lambda_{\max}$ plateaus below $2/\eta$. Smaller batch sizes lead to lower plateau values.
In the context of mini-batch algorithms, (i) Jastrzębski et al. [4, 5] noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when. Moreover, Cohen et al. [1] noticed that SGD somehow acclimates to the hyperparameters. Gilmer et al. [27] established effects of initialization and architecture choices on stability of SGD, without providing a definitive stability condition. These works did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 2, often without a proper progressive sharpening phase. This scenario leaves the most basic questions open: In what way the location of convergence of SGD acclimates to the choice of hyperparameters? What are the key quantities involved? To be more specific, can we characterize the training phenomena in (i), (ii), (iii) above? What determines them? Does SGD train in an unstable regime?
Previous Works on SGD Stability.
A series of works, Wu et al. [28], Ma and Ying [29], Granziol et al. [30], Wu et al. [31], Mulayoff and Michaeli [32], give sharp step-size stability thresholds for constant-step-size SGD on quadratic losses. However, the problem they tackle (is motivated by) but is not if and how SGD shows any EoS-type phenomenon when training neural networks [^1] For instance, in principle the level of stabilization of Frobenius norm for them could be independent of batch size, see Figure 6b in [31], differing from behavior of $\lambda_{\max}$ as described in Figure 2 and by [1, 5, 33]).. Specifically, they leave open (1) what statistics of the Hessians are impacted by progressive sharpening [^2] I.e., in which directions progressive sharpening happens. We show that Batch Sharpness increases—not $\lambda_{\max}$—this implies that the covariance of the Hessians increases at least as fast as the mean; (2) if SGD stabilizes around stationary points or not; (3) in what precise way SGD adapts location of stabilization to hyper parameters, and why. These studies model stability of SGD on quadratics in general, we specify how to characterize the stability given the precise progressive sharpening phenomena of the neural networks landscape. Moreover, from their analysis is non trivial to understand if Batch Sharpness at $2/\eta$ is a possible edge of stability, thus we proved Theorem 1 with a slightly different approach. Empirically, multiple works—e.g., Lee and Jang [17]—establish the presence of oscillations in the SGD trajectory. However, did not distinguish between noise and curvature-driven oscillations—the ones directly relevant to the question of how the curvature adapts to the hyper parameters [^3] As we discuss in Section 3, 5 and Appendix G. .
Flatness and Generalization.
SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance [2, 7, 9, 34, 35, 36, 37, 38]. This advantage has been widely attributed to some notion of flatness of the minima [15, 33, 39, 40, 41, 42, 43]. Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks [44, 45]. Our result is inherently a result about mini-batch training improving flatness. Specifically, we explain why: Training with smaller batches constraints the dynamics to areas with smaller eigenvalues of the full-batch Hessian. This quantifies and characterizes prior observations that SGD tends to locate flat minima and that smaller batch sizes result in reduced Hessian sharpness [2, 33].
(a) Trajectory
(b) Step size schedule
(c) Loss
(d) Zoomed-in
(e) GNI of Eq. (1)
Figure 3: Quadratics: Dynamics of SGD on a 1-D quadratic with $N$ datapoints, $L(x)=\frac{1}{2N}\sum_{i}(x-a_{i})^{2}$, where $a_{i}\sim\mathcal{N}(0,1)$. At step size, oscillations are present. Yet, only when the step size becomes larger than 2/sharpness = 2 (last step size change, red line), the oscillations become unstable (d), the loss diverges (c). Meanwhile, the quantity of (1) consistently stays at $2/\eta$, indicating the presence of Type-1 oscillations.
3 Two Types of Oscillations in SGD
The key defining aspect of EoS is about the solutions found by the algorithm adapting to the optimizer’s hyperparameters. This in the case of full-batch algorithms happens through the appearance of an oscillatory regime. Mini-batch SGD, however, always oscillate because its gradient is noisy and the step size does not vanish. The central question is therefore which oscillations signal curvature-limited dynamics (EoS-like). Both types of oscillation involve quantities stabilizing near the critical threshold of $2/\eta$, yet they differ fundamentally:
  1. 1.
    Type-1 Noise-Driven Oscillation. SGD can wobble around a stationary point simply because gradients vary across batches and the step size is not annealed. This occurs even if the Hessian is small—can arise in flat regions of the landscape—as, fixed–step-size, mini-batch noise alone can induce perpetual “wiggling”. Such noise-driven behavior is well-documented in empirical work [17] and well-studied in classical stochastic approximation [46, 47, 48, 49].
  2. 2.
    Type-2 Curvature-Driven Oscillation. Here, the Hessian’s magnitude and the step size interact so that the algorithm repeatedly “overshoots” the current valley, creating catapult-like surges in loss. In other words, once the local, or perceived, curvature is too large relative to $2/\eta$, the updates become unstable in a manner analogous to the classic EoS [1].
Type-1 oscillations occur even in simple quadratic settings no matter the step size—Figure 3.
Proposition 1.
Denote $L_{B}(\theta)$ the minibatch loss, $L(\theta)$ the full-batch loss, and $\mathcal{H}$ its Hessian, $\mathcal{P}_{b}$ the distribution of the batches of size $b$. Assume the trajectory oscillates (no matter the Type) then
[…]
Crucially, the appearance of some quantity—Gradient-Noise Interaction (GNI)—being $2/\eta$ implies the system is oscillating, not why. It does not mean, in principle, that the landscape or the curvature adapted to the hyper parameters. In the case of Type-1, $2/\eta$ is about the ratio between the covariance of the gradients and the size of the full-batch gradient, independently of the curvature [^4] See Appendix A for the proof of Proposition 1 and a classical treatment of Type-1 oscillations, which characterize any stochastic approximation/otpimization algorithm whenever the step size $\eta_{k}$ is not annealed such that $\sum_{k=1}^{\infty}\eta_{k}<\infty$.. Importantly, a way to differentiate between the two types of oscillation is by changing the hyper parameters and observing the outcome. When the size of oscillations increases (bigger step or smaller batch) the dynamics diverges—catapults away—for Type-2 oscillations; it just increases the size of the oscillations—quickly restabilizing—for Type-1, see Figure 3 and Definition 2 in Section 5.
(a) LR increased early
(b) LR increased late
(c) Batch decreased early
(d) Batch decreased late
[…]